The deepest point in the ocean is in the Mariana Trench, about 11 km deep, in the Pacific. The pressure at this depth is huge, about 1.13 x 10^8 Pa. (a) Calculate the change in volume of 1.00 m3 of seawater carried from the surface to this deepest point. (b) The density of seawater at the surface is 1.03 x 10^3 kg/m3. Find its density at the bottom. (c) Explain whether or when it is a good approximation to think of water as incompressible.l

1
Expert's answer
2021-11-15T12:14:59-0500

Part(a)

d=11km=11000m

"\rho=1.03\times10^3kg/m^3"

"p_1=1.13\times10^8 Pa\\p_2=1.01\times10^5Pa"


"∆P=1.13\times10^{8}-1.01\times10^{5}=1.12899\times10^8Pa"

"\beta=-V\frac{∆p}{∆V}"

"2.1\times10^{9}=-1\times\frac{1.12899\times10^8}{∆V}"

"\u2206V=-0.05377m^3"

Part(b)


"∆\rho=\frac{m∆\rho}{V}=\frac{0.053\times10^3}{1}=0.053\times10^3kg/m^3"

"∆\rho=\rho_2-\rho_1"

"\rho_2=∆\rho+\rho_1"


"\rho_2=(0.053+1.03)\times10^3kg/m^3=1.08\times10^3kg/m^3"

Part(c)

Yes it is a good approximation because in the 11km the density of water changed by only (5-6)%.